OpenStudy (anonymous):

find the equation of the tangent line to (e^xy)+(x^2)-(y^2)=5 through the point (2,0)

OpenStudy (anonymous):

Looks like an implicit differentiation problem to me, I will first have to give it a try myself though.

OpenStudy (anonymous):

\(\large (e^xy)+(x^2)-(y^2)=5 \) OR \(\large (e^{xy})+(x^2)-(y^2)=5 \) and i agree... implicit diffn....

OpenStudy (amistre64):

\[y=\frac{dy}{dx}(x-2)\]

OpenStudy (anonymous):

dpalnc it is the second equation you wrote

OpenStudy (anonymous):

for the derivative i got: (-2x+(e^xy)y)/(e^xy)x-2y

OpenStudy (anonymous):

did anyone get that

OpenStudy (anonymous):

If I use implicit differentiation for the 2nd equation that @dpalnc posted I get \[ \Large e^{xy}\left(y+x\frac{dy}{dx}\right) +2x -2y \frac{dy}{dx}=0\]

OpenStudy (anonymous):

Will have to rearrange that a bit @mariaad

OpenStudy (anonymous):

i got that too but i thought i had to simplify it

OpenStudy (anonymous):

You have to solve for \(\Large \frac{dy}{dx} \)

OpenStudy (anonymous):

to find the slope i would have to plug in 2 to that? because it is the derivative?

OpenStudy (anonymous):

well first solve for dy/dx, then you can do that. because dy/dx=f'(x) and that is the slope at any given point.

OpenStudy (anonymous):

\[ \Large \frac{dy}{dx}=\frac{-2x-ye^{xy}}{xe^{xy}-2y} \] That is what I get after solving for dy/dx, I recommend you to double check though.

OpenStudy (anonymous):

yeahh cus i basically got the same thing except on the nominator i got -2x+ye^xy instead of minus

OpenStudy (anonymous):

so now i plug in 2 to that? but i get (-4+e

OpenStudy (anonymous):

i got -2..is that what you got?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

well yes seems correct to me.

OpenStudy (anonymous):

okk thank you then for the equation i got y=-2x+4

OpenStudy (anonymous):

best thing you could do is feed wolframalpha now with your equations and see if it's correct, if you want to make sure that there is no mistake.

OpenStudy (anonymous):

okk thanks again

OpenStudy (anonymous):

welcome

OpenStudy (amistre64):

just a note: \[\Large e^{xy}\left(y+x\frac{dy}{dx}\right) +2x -2y \frac{dy}{dx}=0\] you will still need to rearrange to find dy/dx; but at this stage you can input the point values to declutter it \[\Large e^{2*0}\left(0+2\frac{dy}{dx}\right) +2(2) -2(0) \frac{dy}{dx}=0\] \[\Large 2\frac{dy}{dx} +4=0\]